The electric potential V as a function of distance x (in meter) is given by: V = (15x2+10x-9) V. The value of electric field of x=1m would be-

Grade:12

4 Answers

Aman Bansal
11 years ago

Dear Varna,

Since field is given as E = \begin{displaymath} E_x = - \frac{dV}{dx}. \end{displaymath}

 

Thus applying this forlumla na dthe substituting x = 1 , we get the field value.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

https://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

chennabasava s patil
11 years ago

16

Varna Sambhangi
11 years ago

i think its 40...hw did you get it as 16??plzz xplain..

Yogita Bang
11 years ago

E = dV/dx = d(15x2 + 10x - 9)/dx = 30x + 10 - 0

Put x=1,

E = 30 + 10 = 40   Smile